OpenStudy (anonymous):

-3(x + n) = x solve for x show steps please!

OpenStudy (anonymous):

-3(x + n) = x -3x -3n = x -3n = 3x + x -3n = 4x -3/4n = x

OpenStudy (anonymous):

thx @Ryan_Kochan , i was really confused on this